PLEASE SHOW WORK FOR ALL!!!

Oxygen Solubility = 0.03 mL O2 /mm Hg PO2/L blood

1.34 mL O2/gram of hemoglobin

Maximal Exercise:

Arterial PO2 (PaO2) = 100 mmHg (~98% saturated)

Venous PO2 (PvO2) = 35 mmHg (~30% saturated)

[Hb] = 130.0 g/L

HR = 190 bpm

Cardiac Output = Q = 20.0 L/min

Use the above information to answer #1-9 ^^

1. What is the arterial oxygen content bound to hemoglobin? _________ mL O2/L blood

2. What is the arterial oxygen content dissolved in blood? _________ mL O2/L blood

3. What is the total arterial oxygen content (CaO2)? _________ mL O2/L blood

4. What is the venous oxygen content bound to hemoglobin? _________ mL O2/L blood

5. What is the venous oxygen content dissolved in blood? _________ mL O2/L blood

6. What is the total venous oxygen content (CvO2)? _________ mL O2/L blood

7. What is the a-vO2 difference? _________ mL O2/L blood

8. What is the VO2? _________ L O2/min (note the units)

9. What is the stroke volume? _________ mL/beat (note the units)
Answer #1

Ans:- 1) arterial O2 containt= ) 1.34 gm O2 in 1gram of Hb

And 130 gm Hb is present in 1lit blood (given)

So arterial containt of O2 = 1.34 X 130 = 174.2 ml O2/ L blood

2) 0.03 ml O2 is dissolved in 1lit of blood at 1mm of HG so

At 100 mm of Hg 0.03 X 100 = 3 ml of O2 / lit of blood

3) Total arterial O2 containt = O2 present in Hb + O2 dissolved in plasma = 174.2 + 3 = 177.2 ml O2 / L of blood

4) as O2 saturation in artery is 174 which is 98 % O2 which can be approx taken as 100% saturation. As venous blood is only 30% saturation so in venous blood amount of O2 bound into Hb = 174 X 30/100 = 174x 0.3= 52.2 ml O2/ L blood